# Shortcut Learning in In-Context Learning: A Survey

### v1

@@ 3.2 Skewed demonstrations
Corresponding to LLMs training, skewed demonstrations are primarily responsible for the generation of acquired shortcuts, as LLMs learn from and reason based on the models presented in the demonstrations. If the demonstrations are skewed, the corresponding reasoning results will also be affected. Many studies have confirmed this by adding specific shortcut triggers to the demonstrations, where the shortcuts in the demonstrations caused a decline in the performance of LLMs Tang et al. (2023); Si et al. (2023b); Levy et al. (2023).
@@ 3.3 LLMs size
While larger LLMs exhibit exceptional capabilities in semantic understanding and reasoning, numerous studies have shown that increasing the LLMs’ scale does not mitigate shortcut learning Kang and Choi (2023); Zhou et al. (2024d); Sclar et al. (2024). Rather, even larger LLMs are more prone to exploiting shortcuts, whether they are instinctive shortcuts or acquired shortcuts Razeghi et al. (2022); Tang et al. (2023); Balepur et al. (2024); Yuan et al. (2024); Han et al. (2024). Although not directly stated, Pan et al. (2023) provide some explanation for this phenomenon: larger LLMs can utilize more demonstrations to enhance their ability to learn new input-label mappings, whereas smaller LLMs cannot. Consequently, when there are more shortcuts in the context, larger LLMs are more susceptible to being influenced by these shortcuts.
@@ 4 Benchmarks of Shortcut Learning
ICL can encompass many of NLU tasks, targeting different types of shortcuts. Therefore, we deem it necessary to show existing evaluation benchmarks to support related research. Specifically, we summarize the datasets used in existing research, along with the corresponding quantitative metrics.
@@ 4.1 Datasets
Most studies have not developed shortcut evaluation datasets, but instead utilize existing datasets Zhao et al. (2021); Fei et al. (2023); Gupta et al. (2023) or make simple modifications to adapt them to their own shortcut evaluation tasks. The most common approach is to add or delete elements from the samples in existing datasets to construct samples that contain shortcuts. For exampple, Tang et al. (2023); Si et al. (2023b) introduce shortcuts by injecting shortcut trigger words, typically a sequence of characters, into the samples, and used this method to evaluate the impact of these shortcuts on ICL. Alternatively, shuffling or deleting option IDs can be used to evaluate the impact of option order on ICL Zheng et al. (2024). Some methods also select a subset of samples that meet certain criteria from existing datasets for experimentation. For example, Sun et al. (2024) choose samples containing words like ‘film’ or ‘movie’ for lexicon shortcuts. LLMs can also play a role for dataset modifications, such as extracting concepts from text for concept shortcuts Zhou et al. (2024c).
There are also methods that propose corresponding benchmarks for different types of shortcuts to better evaluate them, as the the datasets mentioned above are not specifically designed for shortcut evaluation. Shortcut Maze is a comprehensive benchmark for text classification that categorizes shortcuts into lexicon, style, and concept Zhou et al. (2024d). Similarly, Shortcut Suite includes lexicon, overlap, position, and style shortcuts, but it focuses its research on NLI tasks Yuan et al. (2024). Apart from classification tasks, QA benchmarks are also a common scenario for exploring shortcuts. ShortcutQA performs counterfactual editing of samples by adding or excluding shortcut triggers based on the confidence of the target model to obtain shortcut-related samples. It involves various types of shortcuts, including entity, overlap, position, and more Levy et al. (2023). EUREQA is a complex multi-hop QA dataset designed to reduce semantic relevance by adding deceptive semantic cues that distract attention and are unrelated to the correct answer, and to measure the ability of LLMs to perform extensive chained reasoning processes Li et al. (2024a). MMLU-Pro+ introduces various interfering shortcuts into the correct options of reasoning questions across 14 different domains to enhance its ability to evaluate high-order reasoning skills in LLMs Taghanaki et al. (2024). ReWild focuses more on testing LLMs at a fine-grained level in various aspects such as mathematics, logic, and common sense, including the selection and execution of strategies, as well as the tendency to adopt undesirable shortcuts Yang et al. (2024c). ConvRe systematically evaluates the ability of LLM to recognize and process inverse relations, determining whether these inverse relations are affected by shortcuts through the mutual matching of knowledge triples with their corresponding textual descriptions Qi et al. (2023).
@@ 4.2 Metrics
The most commonly used metrics in shortcut evaluation are general evaluation methods for NLP tasks such as accuracy, AUC score Chen et al. (2023) as well as F1 score in classification tasks Tang et al. (2023); Fei et al. (2023), Hits@n in retrieval related tasks Kang and Choi (2023); Pezeshkpour and Hruschka (2024), and exact match score in machine reading comprehension tasks Zhang et al. (2024a); Wang et al. (2023a). If the degree of shortcut learning in ICL is significant, then the corresponding perfomance will also undergo changes Zhou et al. (2024c); Han et al. (2024); Zhou et al. (2024d). Similarly, Fluctuation Rate is designed to represent the ratio of the change in prediction results to the total after a sample is perturbed Wei et al. (2024). In other studies, Conflict Rate is also equivalent to Fluctuation Rate Zheng et al. (2024). Taghanaki et al. (2024) define the Shortcut Selection Ratio, which is used to quantify the tendency of LLMs to maintain their original choices after introducing distracting items.
@@ 5 Mitigation of Shortcut Learning
Given the widespread existence and dangers of shortcuts, numerous studies have investigated corresponding methods to mitigate them. Based on the differences in the targets addressed by these mitigation methods, we categorize the existing approaches into three types: data-centric, model-centric, and prompt-centric.
@@ 5.1 Data-centric Approach
As summarized in Section 3, the shortcuts in LLMs are largely related to the distribution of training data. Therefore, an intuitive approach is to sample task-relevant examples to obtain balanced training data. Kang and Choi (2023) alleviate the dependence on shortcuts by filtering out samples with high co-occurrence probabilities and guiding the retraining of LLMs. Zhou et al. (2024c) employ counterfactual samples generated by LLMs for upsampling and re-trains to mitigate the reliance on shortcuts. Nakada et al. (2024) proposes a systematic oversampling method that leverages the capabilities of LLMs to generate high-quality synthetic data for minority groups. However, the cost of fine-tuning LLMs is too high, and it can also result in catastrophic forgetting Li and Lee (2024). Therefore, there are relatively few corresponding studies.
@@ 5.2 Model-centric Approach
Model pruning alleviates the impact of shortcuts on the model itself by removing biased neurons. This operation can be supported by research on the relationship between neurons and knowledge storage, where neurons are the basic units for storing knowledge and patterns in LLMs Zhao et al. (2024a). In this process, some commonly used post-hoc interpretable methods, such as Integrated Gradients Sundararajan et al. (2017), are used to locate the neurons activated during shortcut learning. By removing these neurons, LLMs can be prompted to explore alternative and correct reasoning paths Ju et al. (2024); Yang et al. (2024b). Generally, the weights of neurons with high shortcut scores are set to 0 to achieve the purpose of pruning Ali et al. (2024). By quantifying the contribution of feedforward neural networks and attention heads to prediction results, Zhou et al. (2024b) identify biased components of LLMs and alleviates their prompt brittleness by removing the corresponding components. However, the negative impact of such component removal on LLMs is still debatable, as neurons have been observed to be polysemous, meaning that a single neuron can be activated by multiple unrelated terms Liu et al. (2022); Bills et al. (2023). That means when shortcut components are removed, some corresponding useful knowledge may also be discarded, introducing new noise.
@@
Calibration corrects biased outputs of LLMs by modifying theirs probability distribution Guo et al. (2017). Contextual Calibration obtains context-driven biased distributions by replacing the sample to be predicted with meaningless inputs, such as ‘N/A’ Zhao et al. (2021). In an ideal unbiased environment, even if the sample to be predicted is set to ‘N/A’, LLMs (Large Language Models) should yield an output similar to 50% Positive and 50% Negative. If the prediction distribution shifts, then this shift is attributed to the corresponding context. Subsequently, during actual predictions, LLMs’ outputs are adjusted based on the biased distribution to mitigate biases in the context.
@@
As research progresses gradually, Contextual Calibration has spawned various calibration methods tailored for different shortcut mitigation scenarios. Prototypical Calibration employs a Gaussian mixture distribution to estimate prototypical clusters for all categories and then utilizes the likelihood of these prototype clusters to calibrate LLMs’ predictions Han et al. (2022). Domain-context Calibration employs randomly selected in-domain words from the task corpus to estimate the label bias of the language model and utilizes this estimation for bias mitigation Fei et al. (2023). Similarly, PMIDC re-weighs scores based on the likelihood of hypotheses (answers) given premises (questions) in a specific task domain Holtzman et al. (2021).
@@
But these calibration schemes may fail due to their inability to effectively estimate contextual bias with only content-free and in-domain random tokens. To address this, Batch Calibration updates the bias as more batches of input data are processed, allowing the bias to stabilize after multiple mini-batches are considered Zhou et al. (2024a). Generative Calibration adjusts the margin of the labels estimated via Monte-Carlo sampling over the in-context model to simply calibrate the contextual prediction distribution Jiang et al. (2023). In-Context Calibration calibrates the test distribution of the demonstration using the prior semantics expected from the samples Jang et al. (2024). PriDe is applied to the task of multiple choice questions by arranging the option content on a small number of test samples to estimate prior information, and then using the estimated prior information to debias the remaining samples Zheng et al. (2024). Unlike the aforementioned probability-based calibration methods, NOISYICL calibrates language models by adding Gaussian noise to the parameters of LLMs, thereby reducing the prediction bias and unreliable confidence in ICL Zhao et al. (2024b).
@@ 5.3 Prompt-centric Approach
Shortcut-based method mitigates the unhealthy focus of LLMs on shortcuts by masking or perturbing shortcut words or concepts. Zhou et al. (2024c) masks words with high relevance to shortcut concepts, where relevance is measured using Pointwise Mutual Information. Wang et al. (2023a) replace entities that may rely on shortcuts with placeholders and alleviates the focus on a particular entity by guiding LLMs to understand the placeholders as a set of similar entities.
@@
Instruction format-based method guides LLMs to produce unbiased responses by modifying the format of the prompt or adding more detailed instructions. For example, it is advisable to utilize semantic verbalizers as much as possible to ensure that the label tokens have semantic relevance to the task Si et al. (2023a). Alternatively, in multiple-choice scenarios, the multiple options can be rearranged, and the rearranged results can be integrated using a majority voting method from LLMs to mitigate the risks associated with shortcuts related to option order Pezeshkpour and Hruschka (2024); Wang et al. (2024). Sun et al. (2024) recommend avoiding overloading the LLMs with too much task information that could lead to remembering previous false associations by mixing prompts and labels. Mixing prompts is achieved by substituting the first sample in the demonstration for mixed prompts, while mixed labels involves randomly replacing the labels with meaningless symbolic tags. In addition to the aforementioned modifications to prompt formats, some methods also use hints Qi et al. (2023) or Chain-of-Thought (CoT) Yuan et al. (2024) to guide LLMs in making more reasonable inferences. Zhang et al. (2024a) utilize the chain of thought generated by LLMs as a mediator variable, and adjust the causal relationship between the input prompt and the output answer through front-door adjustment to mitigate ICL bias.
@@
Prompt search-based method aims to mitigate biases acquired by LLMs from context by searching for unbiased prompts. Unlike instruction format-based method, this approach focuses on retrieving reasonable prompts within a certain range, rather than having a clear modification target. Prompt search-based method can easily be associated with some recent retrieval-augmented approaches, which enhance the quality of ICL by retrieving samples from the corpus that are helpful for predictions Yoran et al. (2024); Ram et al. (2023). Chen et al. (2023) observe a strong negative correlation between the sensitivity and accuracy of ICL, indicating that predictions sensitive to perturbations are unlikely to be correct. Based on this, SENSEL is proposed to avoid providing users with incorrect predictions by discarding samples that are sensitive to perturbations. Lu et al. (2022) utilize the predicted label distribution statistics and propose an entropy-based metric to measure the quality of candidate prompts with diffenret order. Gonen et al. (2023) assumes that lower perplexity correlates with better performance and selects prompts with low perplexity as input.

### v2

@@ 3.2 Skewed Demonstrations
Corresponding to LLMs training, skewed demonstrations are primarily responsible for the generation of acquired shortcuts, as LLMs learn from and reason based on the models presented in the demonstrations. If the demonstrations are skewed, the corresponding reasoning results will also be affected. Many studies have confirmed this by adding specific shortcut triggers to the demonstrations, where the shortcuts in the demonstrations caused a decline in the performance of LLMs [55, 56, 61].
@@ 3.3 LLMs Size
While larger LLMs exhibit exceptional capabilities in semantic understanding and reasoning, numerous studies have shown that increasing the LLMs’ scale does not mitigate shortcut learning [17, 59, 39]. Rather, even larger LLMs are more prone to exploiting shortcuts, whether they are instinctive shortcuts or acquired shortcuts [42, 55, 44, 58, 64]. Although not directly stated, Pan et al. provide some explanation for this phenomenon: larger LLMs can utilize more demonstrations to enhance their ability to learn new input-label mappings, whereas smaller LLMs can not [35]. Consequently, when there are more shortcuts in the context, larger LLMs are more susceptible to being influenced by these shortcuts.
@@ 4 Benchmarks of Shortcut Learning
ICL can encompass many of NLP tasks, targeting different types of shortcuts. Therefore, we deem it necessary to show existing tasks and evaluation benchmarks to support related research on different situations. Specifically, we summarize the the NLP tasks related to shortcut learning, the datasets used in existing research, along with the corresponding quantitative metrics.
| Text Classification | Information Extraction | Natural Language Inference | Cloze | QA | Reasoning | |
| Vanilla-label Bias | [32, 33, 35, 37] | [32, 36] | [35, 37] | [32, 17] | [33, 36, 14] | [34] |
| Context-label Bias | [32, 38, 39] | [32] | ✓ | [32] | [16, 41, 39] | ✓ |
| Domain-label Bias | [43] | [45] | [15, 43] | ✓ | [43, 44] | [42, 34] |
| Reasoning-label Bias | [47, 50, 48, 52, 79, 53, 49, 51, 54] | |||||
| Lexicon | [55, 56, 59, 60] | [55] | [56, 58, 57, 60] | ✓ | [60] | [60] |
| Concept | [12, 59] | ✓ | ✓ | ✓ | ✓ | ✓ |
| Overlap | [56, 58, 57] | [61] | ||||
| Position | ✓ | ✓ | [58] | ✓ | [61, 81] | [62] |
| Text Style | [55, 59] | ✓ | [58] | ✓ | ✓ | ✓ |
| Group Dynamics | [32, 63] | [32] | [63] | ✓ | [63, 64, 44] | ✓ |
4.1 Tasks
We give five common NLP tasks that explore shortcut learning and related research in Table 1. Due to the intersection between different NLP tasks, we briefly state the tasks presented in Table 1.
Information Extraction aims to extract predefined information from unstructured text, typically tasks including relations, entities, and events [45]. In some cases, relation extraction may degenerate into an relation classification task [88]. But in our taxonomy, information extraction is a multivariate task, such as relation extraction, which needs to obtain both entity pairs and triples of their relations.
Natural Language Inference is a typical task that involves two text branches, which determines whether a statement is relevant to the premise and logically reasonable given the premise. It determines whether the second sentence is an Entailment, Contradiction, or Neutral of the first sentence by judging the relationship between two sentences [58].
Cloze is a context-based prediction task that typically involves removing one or more words from a sentence or paragraph, and then asking the model to predict these removed words based on the context [89].
QA. Research in the field of QA is large and diverse, and QA can be classified into different categories based on factors such as context type, context source, type of reasoning, answer format, and so on [90]. In terms of task form, reasoning is sometimes presented in the form of QA. But in the QA tasks discussed in this survey, complex reasoning is not required, rather, the focus is on retrieving or selecting relevant information from the text and generating answers. The emphasis of the question answering task is on information retrieval and option selection, not on deriving new conclusions, such as in multiple-choice QA tasks that consider shortcuts like prompt format, option order, etc. [38], or in machine reading comprehension tasks that involve overlap shortcuts [61].
Reasoning is a process of integrating multiple types of knowledge to draw new conclusions about the world [91], which challenges LLMs’ ability to perform mathematical reasoning [34], causal reasoning [52], multi-hop reasoning, and other complex cognitive processes [49]. It is important to note that, in some cases, NLI is considered a natural language understanding task [92], while in other cases, it is viewed as a reasoning task [93]. In our survey, NLI is excluded from reasoning.
Furthermore, we observe the relationship between shortcut types and tasks. Firstly, Reasoning-label Bias is highly coupled with reasoning tasks, which is why it does not manifest in other types of tasks. Similarly, Overlap only appears in tasks that contain different branches, making it mutually exclusive with text classification, information extraction, and cloze tests. In the context of reasoning, if the text contains only one branch, overlap is not applicable, such as in the derivation of arithmetic sequences. Conversely, if the task involves multiple branches, where overlapping content exists between different branches, like in multi-hop question answering, then overlap may have an impact. Additionally, we have also observed the absence of certain shortcut types in specific tasks, such as the impact of Text Style on the cloze ability of LLMs. For certain rare text styles, the cloze ability of LLMs may decline.
@@ 4.2 Datasets
Most studies have not developed shortcut evaluation datasets, but instead utilize existing datasets [32, 15, 63] or make simple modifications to adapt them to their own shortcut evaluation tasks. The modification relies on researchers’ prior knowledge of shortcuts. The most common approach is to add or delete elements from the samples in existing datasets to construct samples that contain shortcuts. For exampple, studies introduce shortcuts by injecting shortcut trigger words, typically a sequence of characters, into the samples, and use this method to evaluate the impact of these shortcuts on ICL [55, 56]. In addition to simple tokens or n-grams, context-free sentences can also be used as triggers [58]. Alternatively, shuffling or deleting option IDs can be used to evaluate the impact of option order on ICL [14]. Some methods also select a subset of samples that meet certain criteria from existing datasets for experimentation. For example, Sun et al. choose samples containing words like ‘film’ or ‘movie’ for lexicon shortcuts [57]. LLMs can also play a role for dataset modifications, such as extracting concepts from text for concept shortcuts [12].
There are also methods that propose corresponding benchmarks for different types of shortcuts to better evaluate them, as the the datasets mentioned above are not specifically designed for shortcut evaluation. Shortcut Maze is a comprehensive benchmark for text classification that categorizes shortcuts into lexicon, style, and concept [59]. Similarly, Shortcut Suite includes lexicon, overlap, position, and style shortcuts, but it focuses the research on NLI tasks [58]. Apart from classification tasks, QA benchmarks are also a common scenario for exploring shortcuts. ShortcutQA performs counterfactual editing of samples by adding or excluding shortcut triggers based on the confidence of the target model to obtain shortcut-related samples. It involves various types of shortcuts, including entity, overlap, position, and more [61]. EUREQA is a complex multi-hop QA dataset designed to reduce semantic relevance by adding deceptive semantic cues that distract attention and are unrelated to the correct answer, and to measure the ability of LLMs to perform extensive chained reasoning processes [49]. MMLU-Pro+ introduces various interfering shortcuts into the correct options of reasoning questions across 14 different domains to enhance its ability to evaluate high-order reasoning skills in LLMs [68]. ReWild focuses more on testing LLMs at a fine-grained level in various aspects such as mathematics, logic, and common sense, including the selection and execution of strategies, as well as the tendency to adopt undesirable shortcuts [47]. ConvRe systematically evaluates the ability of LLM to recognize and process inverse relations, determining whether these inverse relations are affected by shortcuts through the mutual matching of knowledge triples with their corresponding textual descriptions [66].
@@ 4.3 Metrics
The most commonly used metrics in shortcut evaluation are general evaluation methods for NLP tasks such as accuracy, AUC score [69] as well as F1 score in classification tasks [55, 15], Hits@n in retrieval related tasks [17, 41], and exact match score in MRC tasks [51, 36]. If the degree of shortcut learning in ICL is significant, then the corresponding perfomance will also undergo changes [12, 64, 59]. Similarly, Fluctuation Rate is designed to represent the ratio of the change in prediction results to the total after a sample is perturbed [16]. In other studies, Conflict Rate is also equivalent to Fluctuation Rate [14]. Asgari et al. define the Shortcut Selection Ratio, which is used to quantify the tendency of LLMs to maintain their original choices after introducing distracting items [68]. Overall, the evaluation criteria for shortcuts are closely tied to the specific tasks they belong to, with task-related changes in performance serving as the primary basis for assessment.
@@ 5 Mitigation of Shortcut Learning
Given the widespread existence and dangers of shortcuts, numerous studies have investigated corresponding methods to mitigate them. Based on the differences in the targets addressed by these mitigation methods, we categorize the existing approaches into three types: data-centric, model-centric, and prompt-centric as shown in Figure 7.
@@ 5.1 Data-centric Approach
As summarized in Section 3, shortcuts in LLMs are largely related to the distribution of training corpus. Therefore, an intuitive approach is to sample task-relevant examples to obtain balanced training data. Kang et al. alleviate the dependence on shortcuts by filtering out samples with high co-occurrence probabilities and guiding the retraining of LLMs [17]. Zhou et al. employ counterfactual samples generated by LLMs for upsampling and re-training to mitigate the reliance on shortcuts [12]. Nakada et al. propose a systematic oversampling method that leverages the capabilities of LLMs to generate high-quality synthetic data for minority groups [70]. However, the cost of fine-tuning LLMs is too high, and it can also result in catastrophic forgetting [94]. Therefore, there are relatively few corresponding studies and focus on LLMs with small parameter numbers.
@@ 5.2 Model-centric Approach
Model pruning alleviates the impact of shortcuts by removing biased neurons. This operation can be supported by research on the relationship between neurons and knowledge storage, where neurons are the basic units for storing knowledge and patterns in LLMs [95]. In this process, some commonly used post-hoc interpretable methods, such as Integrated Gradients [96], are used to locate the neurons activated during shortcut learning. By removing these neurons, LLMs can be prompted to explore alternative and correct reasoning paths [50, 71]. Generally, the weights of neurons with high shortcut scores are set to 0 to achieve pruning [72]. By quantifying the contribution of feedforward neural networks and attention heads to prediction results, some studies identify biased components of LLMs and alleviates their prompt brittleness by removing the corresponding components [73]. However, the negative impact of component removal on LLMs is still debatable, as neurons have been observed to be polysemous, meaning that a single neuron can be activated by multiple terms [97, 98]. So when shortcut components are removed, some corresponding useful knowledge may also be discarded, introducing new noise.
@@
Calibration corrects biased outputs of LLMs by modifying theirs probability distribution [99]. Contextual Calibration obtains context-driven biased distributions by replacing the sample to be predicted with meaningless inputs, such as ‘N/A’ [32]. In an ideal unbiased environment, even if the sample to be predicted is set to ‘N/A’, LLMs should yield an output similar to 50% Positive and 50% Negative. If the prediction distribution shifts, then it can be attributed to the corresponding context. Subsequently, during actual predictions, LLMs’ outputs are adjusted based on the biased distribution to mitigate shortcut learning.
@@
As research progresses gradually, Contextual Calibration has spawned various calibration methods tailored for different shortcut mitigation scenarios. Prototypical Calibration employs a Gaussian mixture distribution to estimate prototypical clusters for all categories and then utilizes the likelihood of these prototype clusters to calibrate LLMs’ predictions [65]. Domain-context Calibration employs randomly selected in-domain words from the task corpus to estimate the label bias of the language model and utilizes this estimation for bias mitigation [15]. Similarly, PMIDC re-weighs scores based on the likelihood of hypotheses (answers) given premises (questions) in a specific task domain [33].
@@
But these calibration schemes may fail due to their inability to effectively estimate contextual bias with only content-free and in-domain random tokens. To address this, Batch Calibration updates the bias as more batches of input data are processed, allowing the bias to stabilize after multiple mini-batches are considered [74]. Generative Calibration adjusts the margin of the labels estimated via Monte-Carlo sampling over the in-context model to simply calibrate the contextual prediction distribution [67]. In-Context Calibration calibrates the test distribution of the demonstration using the prior semantics expected from the samples [75]. PriDe is applied to the task of multiple choice questions by arranging the option content on a small number of test samples to estimate prior information, and then using the estimated prior information to debias the remaining samples [14]. Unlike the aforementioned probability-based calibration methods, NOISYICL calibrates language models by adding Gaussian noise to the parameters of LLMs, thereby reducing the prediction bias and unreliable confidence in ICL [76].
@@ 5.3 Prompt-centric Approach
Shortcut-based method. It mitigates the unhealthy focus of LLMs on shortcuts by masking or perturbing shortcut words or concepts. Zhou et al. mask words with high relevance to shortcut concepts, where relevance is measured using Pointwise Mutual Information [12]. Wang et al. replace entities that may rely on shortcuts with placeholders and alleviates the focus on a particular entity by guiding LLMs to understand the placeholders as a set of similar entities [36].
@@
Instruction format-based method. It guides LLMs to produce unbiased responses by modifying the format of the prompt or adding more detailed instructions. For example, it is advisable to utilize semantic verbalizers as much as possible to ensure that the label tokens have semantic relevance to the task [43]. Alternatively, in multiple-choice scenarios, the multiple options can be rearranged, and the rearranged results can be integrated using a majority voting method from LLMs to mitigate the risks associated with shortcuts related to option order [41, 40]. Sun et al. recommend avoiding overloading the LLMs with too much task information that could lead to remembering previous false associations by mixing prompts and labels. Mixing prompts is achieved by substituting the first sample in the demonstration for mixed prompts, while mixed labels involves randomly replacing the labels with meaningless symbolic tags [57].
In addition to the aforementioned modifications to prompt formats, some methods also use hints [66], Chain-of-Thought (CoT) [58], or causal graph [52] to guide LLMs in making more reasonable inferences. For example, Zhang et al. utilize the chain of thought generated by LLMs as a mediator variable, and adjust the causal relationship between the input prompt and the output answer through front-door adjustment to mitigate ICL bias [51]. Task-specific causal graphs can more reliably elicit the causal reasoning ability of language models and alleviate shortcuts caused by causal order conflicts [52].
@@
Prompt search-based method aims to mitigate biases acquired by LLMs from context by searching for unbiased prompts. Unlike instruction format-based method, this approach focuses on retrieving reasonable prompts within a certain range, rather than having a clear modification target. Prompt search-based method can easily be associated with some recent retrieval-augmented approaches, which enhance the quality of ICL by retrieving samples from the corpus that are helpful for predictions [100, 101]. Chen et al. observe a strong negative correlation between the sensitivity and accuracy of ICL, indicating that predictions sensitive to perturbations are unlikely to be correct [69]. Based on this, SENSEL is proposed to avoid providing users with incorrect predictions by discarding samples that are sensitive to perturbations. Lu et al. utilize the predicted label distribution statistics and propose an entropy-based metric to measure the quality of candidate prompts with diffenret order [38]. Gonen et al. assume that lower perplexity correlates with better performance and selects prompts with low perplexity as input [77].